### v2

@@ Abstract
Collective estimation is a variant of collective decision-making, where agents need to achieve consensus on a continuous quantity in a self-organized fashion via social interactions. A particularly challenging scenario is a fully distributed collective estimation with strongly constrained, dynamical interaction networks, for example, encountered in real physical space. In such cases, agents first need to explore a spatially distributed signal through movement, then reach a consensus while being able to communicate only with nearby neighbors. Collectives face several challenges in achieving precise estimation consensus, particularly due to complex behaviors emerging from the simultaneous evolution of the agents’ opinions and the interaction network. While homophilic networks may facilitate collective estimation in well-connected networks [1], we show that disproportionate interactions with like-minded neighbors lead to the emergence of echo chambers, preventing collective consensus. Our agent-based simulation results confirm that, besides a lack of exposure to attitude-challenging opinions, seeking reaffirming information entraps agents in echo chambers. In a potential solution, agents can break free from the pull of echo chambers. We suggest an additional state where stubborn mobile agents (called Messengers) carry data and connect the disconnected clusters by physically transporting their opinions to other clusters to inform and direct the other agents. However, an agent requires a switching mechanism to determine which state to adopt. We propose a generic, novel approach based on a Dichotomous Markov Process (DMP). We show that a wide range of collective behaviors arise from the DMP. We study a continuum between task specialization with no switching (full-time Messengers), generalization with slow switching (part-time Messengers), and rapid task switching (short-time Messengers). Our results show that stubborn agents can, in various ways, help the collective escape local minima, break the echo chambers, and promote consensus in collective opinion dynamics.
@@ 1 Introduction
Among the many different manifestations of collective intelligence, the wisdom of crowds effect [5, 6, 7] stands out as a great candidate for studying collective computational intelligence [4]. While the core idea–that the average of many imperfect estimations can be remarkably close to the true value–seems straightforward, achieving precise estimations depends on meeting certain criteria [6, 8, 9, 10, 11]. Typically, it assumes global knowledge by individuals, which is often not achievable in fully decentralized settings. However, distributed consensus models, particularly DeGroot-like models [12], implement the wisdom of crowds effect without a centralization assumption [13, 8]. These models have also been used to model opinion dynamics, by providing a mechanism for how opinions spread across the interaction network [14, 15, 16, 17]. Achieving a consensus has shown potential applications in collective estimation scenarios and can be used in a variety of engineering fields ranging from machine learning to networks of sensors and groups of robots [18, 19, 1]. Moreover, studying these behaviors gives insights into similar complex systems, as used in economics, politics, and social sciences [20, 15, 8, 21, 22]. Individual-based models of opinion dynamics usually follow a simple set of rules; Examples are Voter models, bounded confidence, and gossip methods [23, 24, 25]. The underlying network of interaction among agents is an important determinant of collective behavior [25, 26]. For example, we know that the connectivity of the network controls the tradeoff between speed and accuracy of consensus formation in collectives with static interaction networks–stronger connected networks promote faster consensus at the expense of decision accuracy [1, 11].
1.1 Homophily and Echo Chambers
Contrary to the assumption of static networks in early models of opinion dynamics, many systems display dynamic networks, as agents tend to constantly rearrange their connections with their peers, making the collective behavior more complex [27]. These rearrangements can be described by specific rules governing the temporal evolution of networks, with homophily being a prominent example. It corresponds to the disproportional tendency of agents to establish links with like-minded neighbors [28, 29, 30]. By adding a new dynamic to the interactions, and hence, altering how the information flows in the network, homophily changes the resulting collective behavior. One potential consequence of homophily in network dynamics is the emergence of ‘echo chambers’: clusters that are highly homogeneous internally and very heterogeneous compared to each other. The strong effects of homophily and echo chambers on collectives highlight their critical importance; particularly, when these effects lead to the emergence of phenomena that are considered threats to human society [31, 32]. An example is their effect on the spread of misinformation [33, 34], where the homogeneity of information within clusters acts as a driver to the diffusion of misinformation [31, 35, 36]. Ultimately, echo chambers reinforce only the locally dominating perspective, foster confirmation biases, and prevent clusters from receiving attitude-challenging information [35, 37, 33]. Other consequences of homophily include the formation of filter bubbles and network segregation [21, 36]. Despite their adverse effects at the collective level, these echo chambers function for a purpose on the individual level: comforting agents with reaffirmation and protecting them from disagreement [35]. Although the actual adversarial outcome of echo chambers in collective dynamics is arguable [38], they indeed inhibit the flow of information on the network in our scenario where achieving consensus is an objective.
1.2 Scenario: Spatial Embedding and Online Sampling of Information
In this work, we consider the scenario of a spatial embedding of opinion dynamics, encountered in the real-world in groups of mobile agents moving through physical space [39, 1]. Many models have already been developed to describe a mechanistic process of how the co-evolution of homophilic networks and opinion dynamics lead to the emergence of echo chambers and complex collective behavior [40, 41, 42, 43]. Although the opinion-updating models coupled to network dynamics vary greatly in the literature, they all typically assume that the initial opinions of agents come from a random pool of possible (discrete) options. Hence, these models do not focus on the origin of the opinions, but rather on their evolution through social interactions. Therefore, by design, these models do not take into account the influence of information aggregation from environmental sources on the evolution of opinions. Contrarily, we assume that the information agents receive is given by a specific information landscape, for example, emitted by one or multiple information sources, and the agent’s position determines the information available to them, hence shaping their opinion. We already know that the position of agents in the network matters [44], but so does the position within the information landscape. The explicit consideration of information space in our model allows us to account for its effects on shaping opinions. Here, the information landscape is modeled by a distribution of environmental signal in 2D Euclidian space, similar to our previous work [1]. Compared to the previous discrete models of initial opinions in homophilic networks, the information landscape provides continuous opinions with arbitrary possible distributions.
1.3 A Motion Model: Homophily as a Pulling Force
Given our agents are embedded in an information landscape, we require new rules to implement homophily in space. Similarly, we define neighborhood as the local proximity (determined by communication range) of agents in the information landscape. This definition of communication range compares to the thresholds in bounded confidence models of opinion dynamics [24, 45]. In the original models of homophily without spatially constrained networks, homophily could affect links between any two random nodes, regardless of their spatial distance [40]. In a spatial network, however, homophily needs to be redefined to account for spatial locality. Therefore, we propose a potential function based on the dissonance value of agents’ opinions, where homophily acts as a local, gravitational pulling force. Homophilic agents actively move and seek sources of information that optimize this value to a minimal disagreement. This way of modeling homophily is in line with the aforementioned functional purpose of homophily–reducing disagreement [35]. As we showed in our previous work [39], the spatial embedding of information is reflected in the collective pattern formation. In particular, the interplay of individual-level sensing of the information field together with the homophilic interaction enables distributed estimation of the global average of the information landscape and results in a self-organized contour-catching behavior, i.e., aggregation of individuals at the iso-lines of the field value [1] (see Fig.1). However, under certain conditions, such as limited communication range, the spatial network can fragment, leading to the formation of ‘echo chambers.’ In non-spatial networks, a significant parameter in promoting the formation of echo chambers is the limited access of agents to attitude-challenging information [33], i.e., information that deviates from the current belief of the agent. In our model with spatial networks, reducing the communication range has a similar effect. In this paper, we study how a limited communication range leads to the emergence of echo chambers, in our spatial scenario.
@@ 1.3 A Motion Model: Homophily as a Pulling Force
Figure 1: Collective contour capturing emerges as the result of collective opinion dynamics and homophily in space. The traces of agents start with black and become blue (and brighter) over time. Agents converge from the initial random distribution to the mean contour lines of the environment for four different information landscapes. Each plot on the left side corresponds to an information distribution on the right grayscale plot, with the iso-contour lines shown in light yellow and the mean contour line (ground truth) in thick blue line. The video of agents optimizing dissonance function following homophily is available via this link: https://figshare.com/s/8c91526b0270c4df8be5
1.4 Breaking-up Echo Chambers with Messengers
To escape the local optimal trap of echo chambers, agents need to receive diverse opinions flowing into their homogeneous clusters. To address the problem in our specific scenario, we introduce a new concept of ‘Messengers’ agents (see details below). These agents serve effectively as data ferries that connect sub-populations across wider distances beyond the communication range limits. It might be achieved when agents can communicate with neighbors out of their filter bubble, or if they can escape from the pull of homophily. Thus, as a solution, besides the ordinary ‘Exploiter’ state, where agents permanently update their opinions by exploiting the information landscape and interacting with others, we introduce a new ‘Messenger’ state. Messenger agents can freely move in space and share their current information with others, but do not acquire any new information, i.e., they do not integrate the opinion of their neighbors and do not sample the information landscape. We investigate if this approach can increase the effective communication range and restore collective consensus. A challenge, however, is to have a decentralized mechanism for agents to decide whether, when, and for how long agents should remain in the Messenger state. Here, we propose a simple decentralized solution based on the Dichotomous Markov Process (DMP) [46]. We study the effect of DMP parameters on individual and collective behaviors, measured by two properties: the ratio of Messengers, and the switching speed between the two states. We then explore the phase diagram of collective behavior and elaborate on an example of collective behavior at each phase.
2 Results
In this section, first, we review consensus formation in the baseline model with only Exploiter agents and study the emergence of echo chambers in low connectivity regimes. We then investigate the ability to restore consensus via the introduction of an additional state of individuals, which we refer to as Messenger, and a DMP governing the stochastic switching between the Exploiter and the Messenger states. The model details are given in the Methods Sec. 4.1.
2.1 Dissensus in Space: The Emergence of Echo Chambers
Figure 2: Normalized number of clusters and opinion precision error decrease by communication range at the beginning and end of simulations for the baseline setup.
As shown previously in [39], the coupling of the local information aggregation and homophily not only allows agents to estimate the mean value of information distribution in the environment but also leads to an emergent spatial collective pattern formation in the information landscape (see Fig. 1.) We refer to this spatial consensus as collective contour-capturing behavior, i.e., the aggregation of agents at the average iso-lines of the information landscape. The resulting patterns show that in seeking specific information sources, individuals actively move in space, and form communities based on the information they receive. Here, we explore the result of such complex behavior in more detail and show that global convergence relies on a sufficiently large communication range. In settings with limited connectivity, however, global consensus breaks down, and only local consensus can be observed. Due to limited connectivity and homophily, the system becomes trapped into local minima, leading to the formation of clusters that serve as echo chambers, with negative consequences on the collective performance due to the inhibition of information flow across the clusters with dissimilar opinions (see Fig. 2).
To quantify how the consensus, and the resulting contour-capturing behavior, rely on sufficient connectivity, we measured the number of clusters being formed as well as the precision of consensus ($E_{\text{P}}^{\text{S}}$, for further information, see Sec. 4.2) at the end of the experiment. A consensus in space is achieved when agents are precisely on the same contour line. We illustrate these two metrics in Fig. 2. The number of clusters and precision error decrease with the communication range after a critical value ($r_{\text{comm}}\approx 0.3$), highlighting a specific regime where the consensus breaks due to the lack of connectivity. This critical value is the minimum required for the initial network to be fully connected. By comparing the final and initial state of the collective in Fig. 2, we observe that the basic method still improves both the cohesion and precision of the collective. In the next sections, we propose and investigate a solution for increasing effective connectivity by utilizing information-carrying mobile agents, ‘Messengers’, with the switching in and out of the Messenger state for individual agents being governed by the DMP.
2.2 Dichotomous Markov Process as a State Switching Mechanism
DMP is a simple stochastic process with the switching dynamics defined by only two parameters $p_{\text{M}}$ and $p_{\text{E}}$. Fig. 4-a illustrates the temporal dynamics of a single agent’s state. Changing the parameter pair influences two properties of a single agent behavior: the ratio of time each agent spends in either of the two states; and how quickly they switch between the two states. The latter, known as the sojourn time (see Methods Sec. 4.1.4), refers to the time between two consecutive switches. We illustrate the time spent in the Exploiter and Messenger states in Fig. 4-a by $\tau_{\text{M}}$ and $\tau_{\text{E}}$, respectively. At the collective level, changing the two parameters affects the collective average sojourn time ($\tau_{\text{S}}$), and the ratio of Messengers ($m$). These two properties are depicted in the 2D parameter space in Fig. 4-b, and Fig. 4-c.
[…]
2.3 Opinion Consensus with Messengers
The ability to arrive at the consensus for the heterogeneous collective consisting of both Messengers and Exploiters depends on two parameters of the DMP ($p_{\text{M}}$ and $p_{\text{E}}$) governing the switching process. A balanced set of properties ($m$ and $\tau_{\text{S}}$) is required to enable Messengers to propagate information efficiently beyond the limitation of the communication range. This also applies to the Exploiters, which are needed to process the information Messengers carry.
Figure 5: Normalized opinion precision error w.r.t the baseline setup indicating different areas in the parameter space. R1: Too few Messengers (too greedy); R2: Set of regions with a balanced number of Exploiters and Messengers, R2-a (Specialized Exploiters): Specialized slow-switching agents with Exploiters in the majority, R2-b (Enhanced Exploration): Fast-switching agents, R2-c (Generalized Messengers): Majority being fast-switching Messengers; R3: Too many Messengers. The supplementary video of the simulations for a sample of each region is available via this link: https://figshare.com/s/61052656b78288a5fb47.
According to the properties of the DMP (see Sec. 4.1.4), the bottom and left parts of the parameter space correspond to the regions with slow switching of states on average (large $\tau_{\text{S}}$), with agents remaining in at least one of the states for very long times. Given that agents in these regions rarely switch their states (see Fig. 4-b), we can consider them as specialized individuals. This is in contrast to the top-right corner, where generalized agents frequently switch between the two states, corresponding to short sojourn times. The generalized agents experience both states during the course of the experiment. Additionally, if we divide the space off-diagonally into two triangles, with the lower right triangle dominated by Exploiters.
2.3.1 Integration-vs-Information Tradeoff
In Fig. 5, as we go up diagonally from the bottom-right corner to the upper-left corner, we encounter collectives with varying proportions of Messengers. For instance, in region R1, Exploiters constitute the largest share of the population on average resulting in strongly clustered collectives, where agents immediately get trapped in local echo chambers. The opinion consensus performance in this region is deficient due to the prevalence of many greedy Exploiters. We consider an Exploiter agent to integrate and process information according to the DeGroot model for social learning (see Sec. 4.1.) Therefore, R1 is dominated by the integration of the available information. Their greedy and exploitative behavior prevents the flow of new information across clusters. In this setting, the system exhibits the most conservative behavior in terms of mobility and tends to over-exploit in terms of information aggregation. For engineering applications, especially where moving is costly or hazardous, selecting parameters closer to this region might be preferred. Conversely, R3 represents an overly dispersed region where extreme exploration of Messengers and inadequate processing of available information negatively impact performance. In this region, too many Messengers, acting as moving information, are redundant; while there are insufficient Exploiters to integrate new information into the collective. In R3, collectives demonstrate the highest network plasticity, meaning any two random agents are likely to meet each other, irrespective of their opinions. R1 and R3 are extremes of either exploitation or exploration, respectively. The high-performing regions are situated between these two extremes, closer to the off-diagonal line (the blue dashed line). We labeled this balanced region as R2. This region contains all local optima between the blue and red contours confining R1 and R3, respectively. Next, we examine different locally optimal regions within R2, focusing on varying switching timescales.
2.3.2 Generalization-vs-Specialization Tradeoff
2.3.3 Temporal Evolution of Parameter Space
(a) t=1k
(b) t=2k
(c) t=4k
(d) t=10k
(e) t=40k
Figure 6: Time development of parameter space for precision error. Bifurcation of the optimal parameters happens for the top-right side of the parameter regime corresponding to fast dynamics, while the bottom-left corner does not undergo a significant shift, showing a more robust performance by the time limit.
In contrast, the local optimum for generalist Exploiters, located in the top right narrow valley in the precision error (R2-b), emerges only when the system has enough time to process the task. It is important to note that this region is advantageous as it minimizes precision error in both the physical and opinion domains, as we discuss further below in Sec. 2.4. Additionally, we observe a bifurcation in the fast-switching parameter regime. Specifically, the small head of the elongated R2 (see Fig. 6-b) grows over time and eventually branches into two distinct regions (R2-b and R2-c). The gap between the two branches widens over time. Another notable temporal shift occurs in the region with generalist agents and a time-averaged majority of Messengers, the larger upper left branch (R2-c,) which gradually ascends toward higher Messenger ratios. This suggests that in scenarios with higher time budgets, increasing exploration efforts–by assigning more agents to perform random walks–is advantageous. The final aspect to highlight is the continuous improvement in the precision of the optimal regions. The precision gain over time demonstrates the existence of a speed-accuracy tradeoff in this scenario. In contrast, extreme parameter setups, such as those characterized by excessive exploitation (R1) and random diffusion of Messenger (R3), do not exhibit any significant improvement over time.
2.4 Contour Capturing Performance
2.4.1 The Effect of Initial State Distribution of the DMP
We already have established that the temporal properties of the DMP significantly affect collective performance. Another important aspect to consider is the distinction between the DMP transient and stationary behaviors. Up to this point, we have assumed that the DMP begins with the expected analytical ratio of Messengers, thereby initially placing the process in its stationary state. However, any deviation from this stationary state puts the system into a transient phase. The progression towards reaching the stationary state is time-dependent and the transient behavior of the DMP also influences the collective performance. Specifically, we show that the transient behavior is linked to the DMP relaxation time ($\tau_{\text{C}}$)  [46] (see Fig. 10 and Eq. Relaxation Time of the DMP in the Appendix). Next, we will explore how the system’s behavior is affected by its initial condition, particularly focusing on the initial population of Messengers. To show this difference in Fig. 7, we conduct a comparison to a case where the system starts without any Messengers. We know that, given fixed transition rates, the system requires time to attain the expected analytical stationary properties, namely the Messenger population ratio.
Same as the temporal development studied earlier in Fig. 6, by comparing the parameter space at different time steps we observe that systems with different time-scales reach their stationary state with different speeds (see Fig. 11). For example, systems with fast-switching dynamics in the top-right corner reach their stationary states more quickly, which makes them more robust to varying initial conditions within the same time constraints. In contrast, slower systems require more time to gradually approach their stationary states. Consequently, we observe that the horizontal tail of R2 shifts downward, eventually aligning with a shape similar to that depicted in Fig. 6-a. With these observations, we highlight the importance of accounting for transient effects, particularly in scenarios where the properties of the system must dynamically change from one configuration to another. They also provide insights into the robustness of each configuration against variations of initial conditions and selecting parameters from different optimal regions.
2.4.2 Different Information Distributions
Our main focus in this paper is on the specific case where the spatial distribution of the information is radial, i.e., the iso-contour lines are circles. This configuration resembles a range of natural situations where a point-like source isometrically emits a concentration into its surroundings, which may then spread through processes like diffusion or convection. To ensure generality, we also tested our model’s performance in environments with various other distributions. The first row of Fig. 8 presents different information distributions, where agents (depicted as cyan dots) converge to the zero-bias (ground-truth) contour line (indicated by a red dashed line). The introduction of Messenger with tuned DMP parameters shows an improved performance across all four environmental benchmarks.
In the second row of Fig. 8, we present the optimal regions within the parameter space for different information distributions. While the shape of these optimal regions varies slightly depending on the distribution, the general features of the error landscape and the underlying mechanisms discussed in previous subsections remain unchanged. Here, we restricted our analysis to the top-right quarter of the parameter space, where the most non-trivial optimal regions are located. Furthermore, we selected a specific optimal parameter pair (marked with the green dot) to demonstrate the final agent spatial distribution for this particular DMP configuration. This visualization confirms that the chosen parameter set results in a satisfactory spatial consensus across all tested information distributions. Additionally, we also adapted our model to an abstract 1-dimensional environment, with a few modifications, such as the implementation of the random walk (see Appendix). The corresponding results show only negligible qualitative differences compared to the 2D environment.
3 Discussion
The integration of explicit spatial behavior into the collective opinion dynamics model introduces a new dimension to the complexity of the model. We demonstrated this through several components: a) initializing opinions based on spatial distribution, b) defining the interaction network according to spatial local proximity, and c) driving agents’ movement in space through homophily. The interplay of these elements results in complex collective behavior, where homophily in space leads to emergent patterns that trace features of the spatial distribution of information, resulting in the so-called ‘contour capturing behavior.’ This collective pattern formation represents a type of spatial consensus, and we show that it is conditional on having sufficient network connectivity. Our simulations for collectives with low connectivity indicate that the greedy behavior of agents leads to the collective becoming trapped in local optima, a process we associate with the emergence of echo chambers. A unique feature of these echo chambers is their similar shapes, which mirror the spatial information distribution. The spatial patterns constructed by the echo chambers reflect the properties of the information landscape and are a consequence of the spatial dynamics of the system, typically ignored in opinion formation models. Also, our findings indicate that these echo chambers inhibit the flow of information, which prevents the collective’s ability to reach a precise consensus. It is important to clarify that the impact of network connectivity is relative to the size of agents’ distribution. Low connectivity can stem from either a limited communication range or a large initial diversity in opinion space. This observation highlights a potential disadvantage of excessive opinion diversity in a collective when the communication range is limited.
We suggested that a method to overcome the local traps inherent in low-connectivity networks is to indirectly extend the effective communication range. This is particularly relevant in systems comprising mobile agents. By leveraging mobility and employing certain individuals as embodied data carriers, individuals can interact beyond the physical limits of their communication. To this end, we introduce a new behavioral role for agents, the Messenger state. This state allows designated agents to function exclusively as carriers of information. A Messenger is a stubborn agent with a fixed opinion, sharing its opinion while moving randomly in space. These agents can be viewed as mobile “quasi-memory” of agents’ opinions, with their independent random movement promoting broader exploration. We employed the Dichotomous Markov Process (DMP) as a simple, distributed switching mechanism between behaviors. The agents decide randomly at each time step whether they should switch to the other state or stay in their current state. A key advantage of this approach is its decentralization, minimal computation requirements, and scalability.
While only varying the DMP parameters and keeping the other parameters of the model constant has allowed for an in-depth exploration of the system, future research could expand on this by investigating the influence of the other parameters of the model. This includes aspects such as collective density, arena size, or information noise. Also, investigating the one-dimensional version, which resembles classical opinion dynamics models, could provide valuable insights into a range of related issues in the field. An important addition to the analysis of these systems is under dynamic information landscapes, where the collective needs to rearrange itself and adapt to the changing environment. We expect that in such settings, the temporal properties of the system would be highlighted even more.

### v3

@@ Abstract
Collective estimation is a variant of collective decision-making where agents reach consensus on a continuous quantity through social interactions. Achieving precise consensus is complex due to the co-evolution of opinions and the interaction network. While homophilic networks may facilitate estimation in well-connected systems, disproportionate interactions with like-minded neighbors lead to the emergence of echo chambers and prevent consensus. Our agent-based simulations confirm that, besides limited exposure to attitude-challenging opinions, seeking reaffirming information entrap agents in echo chambers. To overcome this, agents can adopt a stubborn state (Messengers) that carry data and connect clusters by physically transporting their opinion. We propose a generic approach based on a Dichotomous Markov Process, which governs probabilistic switching between behavioral states and generates diverse collective behaviors. We study a continuum between task specialization (no switching), to generalization (slow or rapid switching). Messengers help the collective escape local minima, break echo chambers, and promote consensus.
@@ 1 Introduction
Among the many different manifestations of collective intelligence, the wisdom of crowds effect [4, 5, 6] stands out as a great candidate for studying collective computational intelligence [3]. While the core idea–that the average of many imperfect estimations can be remarkably close to the true value–seems straightforward, achieving precise estimations depends on meeting certain criteria [5, 7, 8, 9, 10]. Typically, it assumes global knowledge by individuals, which is often not achievable in fully decentralized settings. However, distributed consensus models, particularly DeGroot-like models [11], implement the wisdom of crowds effect without a centralization assumption [12, 7]. These models have also been used to model opinion dynamics, by providing a mechanism for how opinions spread across the interaction network [13, 14, 15, 16]. Such models find applications in engineering domains like distributed sensing, collective robotics, and machine learning [17, 18, 19, 20], as well as provide insights into social systems for economics, politics, and collective behavior [21, 14, 7, 22, 23]. In this work, we study a specific application of opinion dynamics that leads to collective estimation [20], where agents integrate distributed information to arrive at accurate estimation.
To better understand the underlying interaction rules that drive the group-level behavior, many studies rely on agent-based models. These models capture how simple, local rules, when applied at the level of each agent, can lead to emergent consensus or disagreement at the group level. Common examples for opinion dynamics include the Voter models [24], bounded confidence [25], and gossip methods [26]. Beyond the update rules themselves, the structure of interactions plays a critical role in shaping the outcome of opinion dynamics. The underlying interaction network is an important determinant of collective behavior [26, 27]. In particular, as the network becomes sparser, there exists a sparsity threshold above which the network fragments, making consensus unattainable. For example, in static networks, connectivity governs a tradeoff between the speed and accuracy of consensus formation: stronger connectivity leads to faster convergence but can reduce decision accuracy due to premature agreement [20, 10]. This tradeoff becomes even more pronounced in spatially embedded systems, where agents must physically spread out to access diverse information, but increased dispersion risks breaking the network apart.
Contrary to the assumption of static networks in early models of opinion dynamics, many real-world systems exhibit dynamic networks, as agents tend to constantly rearrange their connections with their peers, making the collective behavior more complex [28]. These adaptive interactions introduce additional complexity into collective behavior, as the network structure evolves along with other system states. A prominent mechanism driving such rewiring is homophily: the disproportional tendency of agents to establish links with like-minded neighbors [29, 30, 31], which in turn reshapes information flow in the network.
One notable consequence of homophily is the emergence of ‘echo chambers’: clusters of agents that are internally homogeneous in their opinions and disconnected from other clusters with different opinions. Echo chambers reinforce only the locally dominating perspective, foster confirmation biases, and prevent exposure to attitude-challenging information [32, 33, 34]. These structures disrupt the flow of diverse information across the network and can lead to collective outcomes beyond simple fragmentation. These effects have been linked to phenomena that are considered threats to public discourse and human society [35, 36, 37]. For example, in the spread of misinformation [34, 38], local information homogeneity acts as a driver to the diffusion of misinformation [35, 32, 39]. Other consequences include the formation of filter bubbles and network segregation [22, 39].
Despite their adverse effects at the collective level, echo chambers function for a purpose on the individual level: comforting agents with reaffirmation and protecting them from disagreement [32]. Although the actual adversarial outcome of echo chambers in collective dynamics is arguable [40], their structural role is clear. In this paper, where reaching consensus on distributed information is a key objective, echo chambers are barriers that inhibit information exchange and reduce the precision of collective estimation. In our collective estimation scenario, the presence of a defined objective allows us to examine the effects of homophily and echo chambers in measurable terms and to derive design principles for mitigation in artificial systems.
In this work, we explicitly consider scenario of spatially embedded opinion dynamics, as encountered in real-world groups of mobile agents moving through physical space [41, 20]. While many mechanistic models have been developed to describe the co-evolution of homophilic networks and opinion dynamics, leading to the emergence of echo chambers [42, 43, 44, 45], they typically focus on the evolution of opinions through social interactions, without considering where those opinions originate. In most of these models, agents’ initial opinions are drawn from a random pool of discrete options. Therefore, by design, these models do not take into account the influence of external sources of information on the evolution of opinions, e.g., information gathered from spatial distributions in the environment. Hence, the results of such non-spatial models cannot explain the spatial patterns exhibited in the systems [37, 46].
Contrarily, other models consider scenarios in which agents are embedded in space and influenced by an external environmental signal, either in a fixed ‘cell’, or grid-like environment [47, 48], or a jump-like influence from external information [49]. We assume that the information agents receive is defined by an information landscape, for example, emitted by one or multiple information sources, and the agent’s position determines the information available to them, hence shaping their opinion. We already know that the position of agents in the network matters [50], but so does the position within the information landscape. Compared to other discrete models of initial opinions in homophilic networks, the information landscape can represent continuous information with arbitrary distributions. This landscape represents external information available in space, not directly determining internal opinions. Agents sample it locally and integrate it with social input to form their opinion.
Given the agents in our model are embedded in an information landscape, we define neighborhood as the local proximity of agents, determined by a finite communication range. While not identical, this definition is conceptually comparable to the thresholds in bounded confidence models of opinion dynamics, where only sufficiently similar agents influence each other [25, 51]. However, here the threshold is implicit, emerging from spatial distance instead of explicit opinion similarity, and is decoupled through its embedding in physical space.
To implement homophily in this setting, we require new rules that considers spatiality of the network. In traditional, non-spatial models, homophily can affect links between any pair of nodes, regardless of their spatial distance [42]. However, in spatial systems, interactions are inherently local and limited by proximity. Therefore, we propose a potential function based on the dissonance value of agents’ opinions, where homophily acts as a local, gravitational pulling force, driving agents to actively move and seek sources of information that minimize this disagreement value. This way of modeling homophily is in line with the aforementioned functional purpose of homophily–reducing disagreement [32]. In abstract models of opinion dynamics, such as bounded confidence models, it is well known that small confidence bounds can result in echo chambers formation. In this work, we raise a new question in spatially embedded setting: does limiting communication range lead to similar fragmentation, or does the added spatial structure change the nature of consensus formation? Here, we extend that framework to explore this question and test the conditions under which spatial fragmentation occurs, and whether simple decentralized mechanisms can mitigate it.
Taken together our approach allows us to study emergent phenomena originating from the interplay between spatial dynamics in an (information) landscape, and the distributed opinion formation process. While we model a specific collective estimation scenario, our explicit consideration of agent movements and interactions in space, with their simultaneous impact on the information input, and the evolution of the network, highlights the functional consequences of spatio-temporal self-organization that is difficult to account for in abstract adaptive network models.
@@
Figure 1: Collective contour capturing emerges as the result of collective opinion dynamics and homophily in space. The traces of agents start with black and become blue (and brighter) over time. Agents converge from the initial random distribution to the mean contour lines of the environment for four different information landscapes. Each plot on the left side corresponds to an information distribution on the right grayscale plot, with the iso-contour lines shown in light yellow and the mean contour line (ground truth) in thick blue line. The video of agents optimizing dissonance function following homophily is available via this link: https://doi.org/10.6084/m9.figshare.26217539.v1 [52]
Once clusters form, the reinforcing dynamics of echo chambers may confine agents to their local clusters, limiting their access to diverse information. Here, we ask whether introducing agents with different interaction rules can help reconnect disconnected subgroups and restore collective consensus. Previous research has shown that stubborn (a.k.a. zealot) agents can regulate the formation of echo chambers [53, 54, 55, 37]. Inspired by this idea, we propose a new role for agents in our spatial setting: Messengers. Compared to the abstract models of stubborn agents, Messengers are active in space. These agents serve effectively as data ferries that connect sub-populations across wider distances beyond the communication range limits. Thus, Messengers stand in contrast to the ordinary ‘Exploiter’ state, in which agents continuously update their opinions by optimizing for both conformity and homophily. Messenger agents, by contrast, move freely in space and share their current opinions with others, but do not acquire any new information, i.e., they neither integrate the opinion of their neighbors nor sample the information landscape.
We investigate if this approach can increase the effective communication range and restore collective consensus. A challenge, however, is to define a decentralized mechanism that determines whether, when, and for how long agents remain in the Messenger state. Here, we propose a simple decentralized solution based on the Dichotomous Markov Process (DMP) [56]. We study the effect of DMP parameters on individual and collective behaviors, measured by two properties: the ratio of Messengers, and the switching speed between the two states. By studying the phase diagram of the collective behavior, we aim to identify the characteristic properties of regimes that support consensus. In particular, we investigate how minimal heterogeneity of roles modulates the system’s ability to break echo chambers.